Solve for x over the real numbers:
x(squared) - x -3=0

Answer by stanbon(57377)

Solve for x over the real numbers:
x(squared) - x -3=0
--------
x^2-x-3 = 0
Use the quadratic formula:
x = [1 +- sqrt(1-4*1*-3)]/(2)
----
x = [1 +- sqrt(13)]/2
---
x = (1/2) + (1/2)(sqrt(13) or x = (1/2)-(1/2)sqrt(13)
===================
Cheers,
Stan H.
==================